OpenStudy (anonymous):

help Pythagorean Theorem The length of one leg of a right triangle is 21 ft and the length of the hypotenuse is 29 ft. What is the length of the other leg?

OpenStudy (swaqqout_kid1):

Eminem fan 4life

OpenStudy (jhannybean):



OpenStudy (jhannybean):

pythagorean theorem states hat for right triangles, you take:\[\sf (\text{leg 1})^2 +(\text{leg 2})^2 =(\text{hypotenuse})^2\]in this case, we're given the hypotenuse and a leg, so rearrange the equation to solve for the missing leg. \[\sf (\text{leg 2})^2 = (\text{hypotenuse})^2 - (\text{leg 1})^2\]\[\sf \text{leg 1} = \sqrt{(\text{hypotenuse})^2 - (\text{leg 1})^2}\]

OpenStudy (jhannybean):

Sorry, meant to write \(\text{leg 2} = \sqrt{(\text{hypotenuse})^2 -(\text{leg 1})^2}\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

50.0 feet 25.0 feet 20.0 feet 35.8 feet

OpenStudy (anonymous):

wich one

OpenStudy (jhannybean):

Did you input your values?

OpenStudy (anonymous):

yes and 29-21= 4

OpenStudy (anonymous):

i mean 8 dur but stil that dosn't equal any of the choices

OpenStudy (jhannybean):

You did not square your values. \[\sqrt{(29)^2-(21)^2}= \sqrt{400}\]